Goldman Sachs wants to estimate the amount of an annual healthcare premium needed for an econometrician as part of a new recruiting program. In a sample of 50 econometricians, they found that the average yearly premium needed is $16,000 with a standard deviation of $3,500.
(a) What is the population mean? What is the best estimate of the population mean?
(b) Develop an 90% confidence interval for the population mean.
(c) Assuming a 99% level of confidence, how large of a sample is required with an acceptable error of $400?
Answer #1

A) Expectation of (sample mean)= population mean. Hence, population mean is $16,000. In large samples, generally over 30 we can use law of large numbers and central limit theorem to conclude that sample mean is the best estimate for population mean. This is because the central limit theorem states that in case of large samples, the distribution tends to a normal distribution where E(sample mean)= population mean

B) At 90% the z value is 1.64.

Lower limit of confidence interval is given by: Sample mean-z(standard error of sample)/root n

Upper limit of confidence interval is given by: Sample mean+z(standard error of sample)/root n

Lower limit: 16,000-1.64*3500/root 50= 16,000-811.75=15188.24

Upper limit: 16,000+811.75=16811.75

CI= (15188.24,16811.75)

C) At 99% z value is 2.58.

Upper limit: 16000+2.58*3500/root 50=17277.03

Lower limit:16000-2.58*3500/root 50= 14722.96

CI=(14722.96,17277.03)

Margin of error is given by:

z value* standard error of sample/ root n

400= 2.58*3500/root n

n= 509.63

Sample size required is approximately 510.